OpenStudy (anonymous):

solve the following system of equations x+8y=3 x=7-8y I need and ordered pair if there is one present

OpenStudy (anonymous):

I can show you how to solve this

OpenStudy (anonymous):

http://www.twiddla.com/solved

OpenStudy (anonymous):

I never could really grasp the concept of the system of equations, I have been miserably struggling with it.

OpenStudy (anonymous):

x+8y=3 x=7-8y x+8y=3 x+8y=7 the graph of these lines IS parallel.....so they dont have a solution